# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 2.6 Discussion
We summarize the main results obtained from the toy study. The correct likelihood function for the problem under consideration is given by eq. (7): neglecting the stochastic nature of the MC templates $t_{ji}$, as done for example in eq. (6), has a sizable impact in terms of under-coverage already with an equal amount of data and simulated events. The “Barlow-Beeston lite” or “MC stat.” approaches treat the uncertain MC predictions as an additional source of statistical noise to the data. This accounts in part for the random fluctuations of MC templates, but is not sufficient to restore the right coverage, as demonstrated by the difference between using the lite vs. full Barlow-Beeston methods. When using the correct likelihood function, the asymptotic properties of maximum-likelihood estimators and of the profile-likelihood ratio are manifestly broken even if the number of events per bin is large enough to justify, for example, the approximation of the NLL as a chi-square function. We found that the asymptotic properties can eventually be recovered by applying a sufficiently large scaling of the data and/or simulation sizes, but we could not find a clear indication of which scale should be considered sufficiently “large”.
Approaches based on the post-fit distribution of the PLR suffer from a related problem: the post-fit nuisance parameters, in particular those related with the fluctuations of MC templates, are never exactly set to their true value. The footprint of the biased MC templates survives in the post-fit distribution of $t_{\mu}$, leading to an artificially augmented sensitivity to $\mu$. We remark that the problem of under-coverage of confidence intervals in the presence of nuisance parameters is well-known already conrad2002 ; acero2024 ; Punzi ; Cranmer . It can be ultimately ascribed to the fact that the standard Neyman’s construction for confidence intervals on one parameter of interest is not uniquely defined when the likelihood depends on multiple nuisance parameters, since some prescription on how to project a larger multi-dimensional space onto a one-dimensional closed interval must be supplied Punzi ; Cranmer . The profile-likelihood approach is just one of the infinitely many ways to achieve such a dimensional reduction. While this matter has been mostly addressed in low-statistics counting experiments and for relatively simple problems, we have shown that the problem can also arise in high-statistics measurements.
@@ 2.6.1 An heuristic confidence interval
The rationale behind eq. (23) is the following. The Hessian uncertainty derived from the full Barlow-Beeston likelihood underestimates the real statistical spread of $\hat{\mu}$ due to the presence of spurious constraints which break the asymptotic approximation. Conversely, the Hessian uncertainty $\bar{\sigma}_{{\rm H}}$ would match the standard deviation of $\hat{\mu}$, if the MC templates were exact. We take it as our staring point. Our intuition is that statistical uncertainties on the MC templates should at least add in quadrature to the Poisson uncertainty of the data, which corresponds to the square root in eq. (23). However, assigning an overall uncertainty to the total MC prediction per bin is not fully correct as it ignores the interplay between the $\mu$ parameter and MC statistical fluctuations pertaining to different templates. In fact, the existence of such an interplay is made evident by the breaking of the Barlow-Beeston lite approximation. The ratio between Barlow-Beeston full and lite Hessian uncertainties, corresponding to the last term on the right-hand side of eq. (23), provides an effective correction factor to account for the increased statistical uncertainty when the MC templates $\theta_{ji}$ are allowed to float in the fit independently from each other.
This heuristic definition ($14^{\rm th}$ row of tables 2 and 3) provides good coverage properties for all values of $\mu_{\rm t}$, as well as for different choices of model configurations. This suggests that $\hat{\sigma}_{\rm heur}$ could be at least used as a proxy for the real statistical spread of the $\hat{\mu}$ estimator in the presence of fluctuations of both the data and the simulated samples, and so it provides a practical test to gauge the goodness of asymptotic confidence intervals. This observation is intriguing and perhaps worth being further investigated, although a rigorous proof of its validity for a generic problem is beyond the scope of this work. We should remark, however, that $\hat{\sigma}_{\rm heur}$ is defined in terms of $\bar{\sigma}_{\rm H}$, which for a realistic problem might not be known a priori. In sec. 3.6 we will address the problem of estimating $\bar{\sigma}_{\rm H}$ when only finite-size MC samples are available.
@@ 3 The general case
The main scope of this section is to characterize in more generality the role played by statistical fluctuations of MC templates in the determination of confidence intervals constructed in the asymptotic approximation. For this purpose, the following approach is followed. We start by assuming that the MC templates are exact. This approximation allows us to derive a set of well-known analytic results for point estimators and confidence intervals based on the asymptotic properties of maximum-likelihood estimators Behnke . The same formalism can also accommodate extensions of the likelihood function that account for finite-size effects in an approximate way, such as the Gauss + MC stat. or Barlow-Beeston lite approaches as discussed in section 2. Then, we will treat statistical fluctuations of MC templates as perturbations on top of the true unknown templates and collect their contributions to asymptotic formulas as an overall bias term, whose sign and magnitude will then be the main subject of study.
@@ 3.1 Methodology
Consider a binned data set represented by independent Poisson-distributed event counts ${\bf y}$ with expectation value parametrized by the model function ${\bf f}(\mu,\boldsymbol{\theta})$, where $\mu$ is the parameter of interest and $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{p})$ is a vector of nuisance parameters, which are assumed to be not constrained by auxiliary measurements. The generalization to the case in which both constrained and unconstrained parameters are present is discussed in section 3.7.
In the limit of large event counts, that is $y_{i}\gg 1$ for $i=1,\ldots,n$, the negative log-likelihood function can be approximated by
@@
where ${\bf V}={\rm diag}({\bf y})$, and the dots stand for an overall constant plus terms that are of higher-order in $(y_{i}-f_{i})$. Equation (24) is the well-known Neyman’s $\chi^{2}$ test-statistic Neyman . We take ${\bf f}$ to be a differentiable function of $(\mu,\boldsymbol{\theta})$ and expand it around a suitable initial point $(\mu_{0},\boldsymbol{\theta}_{0})$:
@@
where ${\bf j}$ is the gradient of ${\bf f}$ with respect to $\mu$, ${\bf J}$ is the $n\times p$ Jacobian matrix of ${\bf f}$ with respect to ${\boldsymbol{\theta}}$, and the dots stand for higher-order terms in the Taylor expansion. While this approximation is exact for linear functions, such as the toy model of section 2, for generic functions it should be iterated around the intermediate minima of eq. (24), until convergence [^4] Clearly, the validity of the linear approximation is model dependent and should be assessed case by case. In particular, model functions that change in a very asymmetric way around the true value of the parameters (e.g. due to non-negligible contributions from the second-order derivatives) should be handled with care.. Hence, we will assume that ${\bf j}$ and ${\bf J}$ are the gradients and Jacobian matrix evaluated at a point in the parameter space which is sufficiently close to the global minimum of the NLL function. Without loss of generality, we take ${\bf J}$ as a full-rank matrix, otherwise the model should be redefined to remove the degeneracy.
@@ 3.1.1 The profile-likelihood ratio
Thanks to eq. (25), the function on the right-hand side of eq. (24) can be minimized by solving a linear system of equations, leading to the well-known linear least-square estimators Cowan . To this aim, we can first put the right-hand side of eq. (24) in the more evocative form
$-2\ln L(\mu,{\boldsymbol{\theta}})=\lVert{\bf r}_{\mu}-{\bf A}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})\rVert^{2},$ (26)
where the newly introduced quantities are defined as
${\bf r}_{\mu}={\bf d}-{\bf b}(\mu-\mu_{0})\;\;\;{\rm and}\;\;\;\left({\bf A},{\bf b},{\bf d}\right)\equiv{\bf V}^{-\frac{1}{2}}\left({\bf J},{\bf j},{\bf y}-{\bf f}_{0}\right),$ (27)
The point $\hat{\boldsymbol{\theta}}_{\mu}$ that minimizes eq. (26) at a fixed value of $\mu$ is given by
The matrix ${\bf W}\equiv{\bf A}^{T}{\bf A}$ is positive-definite, hence invertible, because ${\bf A}$ was assumed to have full rank. By inserting the right-hand side of eq. (28) into eq. (26), the log-likelihood function becomes a quadratic function of $\mu$, which can be again minimized analytically yielding the estimator:
$\hat{\mu}=\mu_{0}+\left({\bf b}^{T}{\bf U}{\bf b}\right)^{-1}{\bf b}^{T}{\bf U}{\bf d},$ (29)
with the matrix ${\bf U}$ defined as
@@
After some straightforward algebraic manipulation, we can finally write the profile log-likelihood function as
where $\hat{\boldsymbol{\theta}}$ is the vector of nuisance parameters that minimizes eq. (24). According to Wilks’ theorem Wilks $t_{\mu}$ is asymptotically distributed as a chi-square with one degree-of-freedom, so that the quadratic form
$S={\bf b}^{T}{\bf U}{\bf b}={\bf b}^{T}{\bf b}-{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}$ (32)
converges to the reciprocal of the variance of $\hat{\mu}$. A consequence of eq. (32) is that $\mu$ cannot be measured (i.e. $\sigma_{\rm H}$ would be virtually infinite) when ${\bf b}\in{\rm Span}\{\bf A\}$ [^5] Physically, this would imply that a point in the NP space can be always found such that the right-hand side of eq. (26) is invariant under a variation of $\mu$.: by definition, in this case one can write ${\bf b}$ as a linear combination of the columns of ${\bf A}$, that is ${\bf b}={\bf A}{\bf w}$ for some vector of weights ${\bf w}$, so that
${\bf U}{\bf b}$ $=({\bf 1}-{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}){\bf A}{\bf w}={\bf A}{\bf w}-{\bf A}({\bf A}^{T}{\bf A})^{-1}({\bf A}^{T}{\bf A}){\bf w}$
$={\bf A}{\bf w}-{\bf A}{\bf w}={\bf 0},$ (33)
which then implies $S=0$.
where ${\bf B}$ is a $n\times(p+1)$ matrix obtained by adding the column ${\bf j}$ on the right of ${\bf J}$, and ${\bf C}=\left({\bf B}^{T}{\bf B}\right)^{-1}$ is the covariance matrix of $(\hat{\boldsymbol{\theta}},\hat{\mu})$. By using the formula for the inverse block matrix, one can easily verify that
Equation (35) implies that
${\sigma}^{2}_{\rm H}=\left[{\bf C}\right]_{p+1,\,p+1}=S^{-1},$ (36)
The matrix ${\bf U}$ introduced in eq. (30) is a $n\times n$ symmetric, positive semi-definite, and idempotent matrix, that is ${\bf U}^{T}{\bf U}={\bf U}$, and so is ${\bf 1}-{\bf U}$. Therefore, we can write it as
@@
where ${\bf u}_{j}$ are linearly independent and unit-norm vectors associated with the $p$ unit eigenvalues of ${\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}$. The ${\bf u}_{j}$ eigenvectors provide an orthonormal basis to ${\rm Span}\{{\bf A}\}$, i.e. the linear space spanned by its columns $\{{\bf a}_{1},\ldots,{\bf a}_{p}\}$. Being associated with identical eigenvalues, the vectors ${\bf u}_{j}$ obtained from the eigen-decomposition of ${\bf 1}-{\bf U}$ are not necessarily orthogonal, but can always be made such by using e.g. the Gram–Schmidt process.
Based on eq. (32), two limiting cases can be identified:
@@
  • the vector ${\bf b}$ has a large projection onto ${\rm Span}\{\bf A\}$, that is
    ${\bf b}=\sum_{j=1}^{p}\left({\bf b}^{T}{\bf u}_{j}\right){\bf u}_{j}+{\bf b}_{\perp},$
    with ${\bf b}_{\perp}\notin{\rm Span}\{{\bf A}\}$ and $\lVert{\bf b}_{\perp}\rVert\ll\lVert{\bf b}\rVert$;
  • the vector ${\bf b}$ is almost orthogonal to ${\rm Span}\{{\bf A}\}$.
@@
We see from eq. (32) that the extent to which a problem can be classified according to either of the two cases can be quantified by the positive-definite metric $\rho_{\mu}$ defined by
$\rho^{2}_{\mu}=\frac{{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}}{{\bf b}^{T}{\bf b}}\leq 1,$ (38)
[…]
$1-\left(\left[{\bf C}\right]_{p+1\,p+1}\cdot\left[{\bf C}^{-1}\right]_{p+1\,p+1}\right)^{-1},$ (39)
as defined for example in Minuit JAMES1975343 . Finally, we notice that $\rho^{2}_{\mu}{\sigma}^{2}_{\rm H}$ is also equal to the difference in quadrature between ${\sigma}^{2}_{\rm H}$ and the variance of $\mu$ when $\boldsymbol{\theta}$ are fixed to their post-fit values. Hence, $\rho_{\mu}$ can be in principle computed also from the group impact ($I_{\boldsymbol{\theta}}$) on $\mu$ from all nuisance parameters Pinto:2023yob ; CMS:2024lrd , namely:
3.2 Impact of MC fluctuations on Hessian uncertainties

### v3

@@ 2.8 Discussion
The main results of the toy study are summarized in Table 5. The correct likelihood function for the problem under consideration is given by eq. (8). Neglecting the stochastic nature of the MC templates $t_{ji}$, as done, for example, in eq. (7), has a sizable impact in terms of both the coverage and the size of confidence intervals already with an equal amount of data and simulated events. The “Barlow-Beeston lite” or “MC stat.” approaches treat the uncertain MC predictions as an additional source of statistical noise to the data. This accounts in part for the random fluctuations of MC templates, but is not sufficient to restore the right coverage or interval size, as demonstrated by the difference between using the lite vs. full Barlow-Beeston methods. When using the correct likelihood function in the Gaussian approximation, the asymptotic properties of maximum-likelihood estimators and of the profile-likelihood ratio are manifestly broken even if the number of events per bin is large enough to justify, for example, the approximation of the NLL as a chi-square function. We found that the asymptotic properties can eventually be recovered by applying a sufficiently large scaling of the data and/or simulation sizes, but we could not find a first-principle indication of which scale can be considered sufficiently “large”.
Approaches based on the post-fit distribution of the profile-likelihood ratio in the Guassian approximation suffer from a related problem: the post-fit nuisance parameters, in particular those related with the fluctuations of MC templates, are never exactly set to their true value. The footprint of the fluctuated MC templates survives in the post-fit distribution of $t_{\mu}$, leading to an artificially augmented sensitivity to $\mu$. We remark that the problem of under-coverage of confidence intervals in the presence of nuisance parameters is well-known already [6, 20, 24, 13]. It can be ultimately ascribed to the fact that the standard Neyman’s construction [21] for confidence intervals on one parameter of interest is not uniquely defined when the likelihood depends on multiple nuisance parameters, since some prescription on how to project a larger multi-dimensional space onto a one-dimensional closed interval must be supplied [24, 13]. Approaches based on the profile-likelihood function are just instances of the infinitely many ways to achieve such a dimensional reduction. While this matter has been mostly addressed in low-statistics counting experiments and for relatively simple problems, we have shown that the problem can also arise in high-statistics measurements.
@@ 2.8.1 An heuristic confidence interval
The rationale behind eq. (26) is the following. The Hessian uncertainty derived from the full Barlow-Beeston likelihood underestimates the real statistical spread of $\hat{\mu}$ due to the presence of spurious constraints which break the asymptotic approximation. Conversely, the Hessian uncertainty $\bar{\sigma}_{{\rm H}}$ would match the standard deviation of $\hat{\mu}$, if the MC templates were exact. We take it as our starting point. Our intuition is that statistical uncertainties on the MC templates should at least add in quadrature to the Poisson uncertainty of the data, which corresponds to the square root in eq. (26). However, assigning an overall uncertainty to the total MC prediction per bin is not fully correct as it ignores the interplay between the $\mu$ parameter and MC statistical fluctuations pertaining to different templates. In fact, the existence of such an interplay is made evident by the breaking of the Barlow-Beeston lite approximation. The ratio between Barlow-Beeston full and lite Hessian uncertainties, corresponding to the last term on the right-hand side of eq. (26), provides an effective correction factor to account for the increased statistical uncertainty when the MC templates $\theta_{ji}$ are allowed to float in the fit independently from each other.
This heuristic definition ($14^{\rm th}$ row of tables 3 and 4) provides good coverage properties for all values of $\mu_{\rm t}$, as well as for different choices of model configurations. This suggests that $\hat{\sigma}_{\rm heur}$ could be at least used as a proxy for the real statistical spread of the $\hat{\mu}$ estimator in the presence of fluctuations of both the data and the simulated samples, and so it provides a practical test to gauge the goodness of asymptotic confidence intervals. This observation is intriguing and perhaps worth being further investigated, although a rigorous proof of its validity for a generic problem is beyond the scope of this work. We should remark, however, that $\hat{\sigma}_{\rm heur}$ is defined in terms of $\bar{\sigma}_{\rm H}$, which for a realistic problem might not be known a priori. In sec. 3.6 we will address the problem of estimating $\bar{\sigma}_{\rm H}$ when only finite-size MC samples are available.
@@
MC stat. CI Nominal $k=40$ $\epsilon=0.12$ $n=20$ $N=2\times 10^{8}$
none Hessian $\times$ $\checkmark$ $\times$ $\times$ $\times$
MC stat. Hessian $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
$B$-lite Hessian $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
$B$ Hessian/PLR $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
$B$ PFC/Barlett $\times$ $\checkmark$ $\times$ $\times$ $\times$
$B$ CH $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
$B$ FC Cheat $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$B$ Heuristic $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 5: Summary of the coverage properties for different model configurations. A thick (cross) indicates that the right coverage, within a relative $\pm 5\%$ tolerance, is (not) attained for both the nominal and alternate data sets. In this table, “BB” is a shorthand for Barlow-Beeston.
@@ 3 The general case
The main scope of this section is to characterize in more generality the role played by statistical fluctuations of MC templates in the determination of confidence intervals constructed in the asymptotic approximation. For this purpose, the following approach is followed.
We start by assuming that the MC templates are exact. This approximation allows us to derive a set of well-known analytic results for point estimators and confidence intervals based on the asymptotic properties of maximum-likelihood estimators [4]. The same formalism can also accommodate extensions of the likelihood function that account for finite-size effects in an approximate way, such as the Gauss + MC stat. or Barlow-Beeston lite method as discussed in section 2, which are in practice equivalent to an overall rescaling of the statistical power of the data. Albeit approximate, these methods are still relevant in light of their diffusion in HEP analyses. We will then treat statistical fluctuations of MC templates as perturbations on top of the true unknown templates and collect their contributions to asymptotic formulas as an overall bias term, whose sign and magnitude will then be the main subject of study.
We remark that our treatment is not directly applicable to the full Barlow-Beeston method, because the introduction of additional nuisance parameters (associated with the individual MC predictions) breaks the quadratic dependence of the NLL on the parameters, which is a crucial assumption in our derivation. For example, in the context of the toy study discussed in section 2, the implementation of the full Barlow-Beeston method has been found to affect the coverage of asymptotic confidence intervals in a significant way, although still insufficient to restore the right coverage. We will come back to the validity of the Barlow-Beeston lite approximation at the end of this section. We can argue, however, that the main results derived in this section are expected to be relevant even for the full Barlow-Beeston method. Indeed, the main source of under-coverage that will be identified by our study is related to the occurrence of low-statistics regimes along particular linear combinations of the data, represented in this context by a finite number of event counts, despite the latter being large in absolute: in such situations, asymptotic formulas for maximum-likelihood estimators are likely to be just not applicable even if the exact likelihood functions were considered.
@@ 3.1 Methodology
Consider a binned data set represented by independent Poisson-distributed event counts $\bf y$} with expectation value parametrized by the functional model ${\bf f}(\mu,\boldsymbol{\theta})$, where $\mu$ is the parameter of interest and $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{p})$ is a vector of nuisance parameters, which are assumed to be not constrained by auxiliary measurements. The generalization to the case where both constrained and unconstrained parameters are present is discussed in C, where an interesting connection between the present work and the main subject of ref. [12] is pointed out.
As discussed previously, in the limit of large event counts, that is, for $y_{i}\gg 1$ for $i=1,\ldots,n$, the negative log-likelihood function can be approximated as
@@
where ${\bf V}={\rm diag}({\bf y})$, and the dots stand for an overall constant plus terms that are of higher-order in $(y_{i}-f_{i})$. Aside from a factor of $1/2$, the right-hand side of eq. (27) coincides the well-known Neyman’s $\chi^{2}$ test-statistic [2]. We take $\bf f$} to be a differentiable function of both $\mu$ and $\boldsymbol{\theta}$ and expand it around a suitable initial point $(\mu_{0},\boldsymbol{\theta}_{0})$:
@@
where $\bf j$} and $\bf J$} are the $n\times 1$ and $n\times p$ blocks of the Jacobian matrix of $\bf f$} with respect to $\mu$ and $\boldsymbol{\theta}$}, respectively, and the dots stand for higher-order terms in the Taylor expansion. While this approximation is exact for linear functions, such as the toy model of section 2, for generic functions it should be iterated around the intermediate minima of eq. (27), until convergence [^4] The validity of the linear approximation is model dependent and should be assessed case by case. In particular, functional models that change in a very asymmetric way around the true value of the parameters (e.g. due to non-negligible contributions from the second-order derivatives) should be handled with care.. Hence, we will assume that $\bf j$} and $\bf J$} are evaluated at a point in the parameter space which is sufficiently close to the global minimum of the NLL function. Without loss of generality, we take $\bf J$} as a full-rank matrix, otherwise the model should be redefined to remove the degeneracy.
@@ 3.1.1 The profile-likelihood ratio
Thanks to eq. (28), the function on the right-hand side of eq. (27) can be minimized by solving a linear system of equations, leading to the well-known linear least-square estimator [10]. To this aim, we can first put the right-hand side of eq. (27) in the more concise form:
$-2\ln L(\mu,{\boldsymbol{\theta}})=\lVert{\bf d}-{\bf b}(\mu-\mu_{0})-{\bf A}(\boldsymbol{\theta}-\boldsymbol{\theta}_{0})\rVert^{2},$ (29)
where $\lVert\cdot\rVert$ indicates the Euclidean norm, while the newly introduced quantities are defined as
${\bf d}$ $={\bf V}^{-\frac{1}{2}}({\bf y}-{\bf f}_{0})$ (30)
${\bf b}$ $={\bf V}^{-\frac{1}{2}}{\bf j}$ (31)
${\bf A}$ $={\bf V}^{-\frac{1}{2}}{\bf J}$ (32)
In the following, we will equivalently refer to $\bf b$} and $\bf A$} as parts of the Jacobian matrix, although they formally differ from $\bf j$} and $\bf J$} for a left multiplication by ${\bf V}^{-\frac{1}{2}}$. As we shall see, the effect of finite-size MC samples on asymptotic confidence intervals can be fully characterized in terms of $\bf b$} and $\bf A$}.
The NP point that minimizes eq. (29) at a fixed value of $\mu$ is provided by the profile likelihood estimator $\hat{\boldsymbol{\theta}}_{\mu}$. By inserting the expression for $\hat{\boldsymbol{\theta}}_{\mu}$ into eq. (29), the log-likelihood function becomes a quadratic function of $\mu$, which can be again minimized yielding the maximum-likelihood estimator $\hat{\mu}$. Explicit formulas for $\hat{\boldsymbol{\theta}}_{\mu}$ and $\hat{\mu}$ are provided in D.1. After some straightforward calculations, we can finally write the profile log-likelihood function [^5] We remark that Eq. (33) would be the exact profile-likelihood ratio only in the Gaussian approximation. as
where the matrix $\bf U$} is defined as
@@
According to Wilks’ theorem [29], $t_{\mu}$ is asymptotically distributed as a chi-square with one degree-of-freedom, so that the quadratic form
$S={\bf b}^{T}{\bf U}{\bf b}={\bf b}^{T}{\bf b}-{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}=\sigma_{\rm H}^{-2}$ (35)
must be equal to the inverse variance of $\hat{\mu}$. We notice that $\sigma_{\rm H}$ can be equivalently obtained from the diagonal elements of the covariance matrix $\bf C$}, defined as the inverse-Hessian matrix of the NLL function, see eq. (17). A proof of this statement, together with an explicit expression of $\bf C$} in terms of $\bf b$} and $\bf A$}, can be found in D.1.
The matrix $\bf U$} introduced in eq. (34) is a $n\times n$ symmetric, positive semi-definite, and idempotent matrix [^6] A symmetric square matrix $\bf U$} is said to be idempotent if ${\bf U}^{T}{\bf U}={\bf U}={\bf U}^{T}$.. The same properties pertain to the matrix ${\bf 1}-{\bf U}$, where $\bf 1$} denotes the $n\times n$ identity matrix. We exploit the fact that an idempotent matrix can be decomposed as the sum of the outer product of its normalized eigenvectors, and rewrite eq. (34) in the equivalent form
@@
where ${\bf u}_{j}$ are linearly independent vectors of unit norm associated with the $p$ eigenvalues of the square matrix ${\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}$, which are all equal to one. It can be easily proved that the ${\bf u}_{j}$ eigenvectors provide an orthonormal basis to the linear space spanned by the $p$ linearly-independent columns ${\bf a}_{1},\ldots,{\bf a}_{p}$ of $\bf A$}. Being associated with degenerate eigenvalues, the vectors ${\bf u}_{j}$ obtained from the eigen-decomposition of ${\bf 1}-{\bf U}$ are not necessarily orthogonal, but can always be made such by using, for example, the Gram-Schmidt process, so they can be taken as mutually orthogonal with no loss of generality.
The formalism developed so far shows that the Hessian uncertainty $\sigma_{\rm H}$ is simply related to the norm of the projection of the vector $\bf b$} onto the space orthogonal to the linear space spanned by the columns of $\bf A$}. For brevity, we will denote the latter by ${\cal V}_{\bf A}$. Given that the ${\bf u}_{j}$ vectors provide a complete basis of ${\cal V}_{\bf A}$, the $\bf b$} vector can be always decomposed as
${\bf b}=\sum_{j=1}^{p}\left({\bf b}^{T}{\bf u}_{j}\right){\bf u}_{j}+{\bf b}_{\perp},$ (37)
where $\bf b$$_{\perp}$ is vector of dimension $n$ orthogonal to ${\cal V}_{\bf A}$, i.e. ${\bf b}^{T}_{\perp}{\bf u}_{j}=0$ for $j=1,\ldots,p$. It is convenient to consider two limiting cases, namely:
  • 1.
@@
    the $\bf b$} vector has a large projection onto ${\cal V}_{\bf A}$, that is $\lVert{\bf b}_{\perp}\rVert\ll\lVert{\bf b}\rVert$;
  • 2.
    the $\bf b$} vector is almost orthogonal to ${\cal V}_{\bf A}$, so that $\lVert{\bf b}_{\perp}\rVert\approx\lVert{\bf b}\rVert$.
@@
The extent to which a given problem can be classified according to either of the two cases can be quantified by the positive-definite metric $\rho_{\mu}$ defined as [^7] It can be proved that $\rho_{\mu}^{2}$ is equal to the so-called global correlation coefficient, which is used, for example, in Minuit [19] for fit diagnostics. See D.1 for further details.
$\rho_{\mu}=\sqrt{\frac{{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}{\bf b}}{{\bf b}^{T}{\bf b}}},$ (38)
which is always less or equal than one, so that the two cases would correspond to either $\rho_{\mu}\to 1\;{\rm or}\;0$, respectively.
3.2 Impact of MC fluctuations on $\sigma_{\rm H}$